A stock has a beta of 1.21 and an expected return of 11.9 percent. A risk-free asset currently earns 3.85 percent.

a. What is the expected return on a portfolio that is equally invested in the two assets? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Expected return             %

b. If a portfolio of the two assets has a beta of .81, what are the portfolio weights? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., 32.1616.)

Weight of the stock
Weight of the risk-free asset

c. If a portfolio of the two assets has an expected return of 11.1 percent, what is its beta? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Beta            

d. If a portfolio of the two assets has a beta of 2.41, what are the portfolio weights? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 4 decimal places, e.g., 32.1616.)

Weight of the stock
Weight of the risk-free asset
Answer #1

part A:

Expected Ret = Weighted Avg ret

Partculars Weight Ret Wtd Ret
Stock 0.5 11.90% 5.95%
Rf Asset 0.5 3.85% 1.93%
Expected Ret 7.88%

Part B:

Portfolio Beta = Weighted in Stock * Stock Beta

Weight in Stock = Portfolio Beta / Stock Beta

= 0.81 / 1.21

= 0.67

Weigt in Risk Free Asset = 1 - 0.67

= 0.33

Part C:

Partculars Weight Ret Wtd Ret
Stock X 11.90% 0.119X
Rf Asset 1-X 3.85% 0.0385 - 0.0385X
Expected Ret 0.0805X + 0.0385

0.0805X + 0.0385 = 0.111

0.0805X = 0.111 - 0.0385

= 0.0725

X = 0.0725/ 0.0805

= 0.90

Portfolio Beta = Weight In Stcok * Stock Beta

= 0.9 * 1.21

= 1.089

Part D:

Weight In stock = Portfolio Beta / Stock Beta

= 2.41 / 1.21

= 2

Weight In stock = 2

Weight is Risk Free Asset = -1